EMD Corporation manufactures two products, Product S and Product W. Product W is of fairly recent origin, having been developed as an attempt to enter a market closely related to that of Product W. Product W is the more complex of the two products, requiring 2 hours of direct labor time per unit to manufacture compared to 1 hour of direct labor time for Product S. Product W is produced on an automated production line.

Overhead is currently assigned to the products on the basis of direct-labor-hours. The company estimated it would incur $1,038,898 in manufacturing overhead costs and produce 14,100 units of Product W and 70,500 units of Product S during the current year. Unit cost for materials and direct labor are:

Product S Product W
Direct material $ 14 $ 29
Direct labor 15 14

Required:

a-1. Compute the predetermined overhead rate under the current method of allocation.

a-2. Determine the unit product cost of each product for the current year.

b. The company's overhead costs can be attributed to four major activities. These activities and the amount of overhead cost attributable to each for the current year are given below:

Total Activity
Activity Cost Pool Total Cost Product S Product W Total
Machine setups required $ 481,000 1,210 2,490 3,700
Purchase orders issued 53,040 543 137 680
Machine-hours required 286,350 7,780 11,310 19,090
Maintenance requests issued 218,508 672 900 1,572
$ 1,038,898

Using the data above and an activity-based costing approach, determine the unit product cost of each product for the current year.
Answer #1

a1) Overhead rate = 1038898/98700 = 10.53 per hour

a2) Unit product cost

Product S Product W
Direct material 14 29
Direct labor 15 14
Overhead 10.53 10.53*2 = 21.06
Unit product cost 39.53 64.06

b) Unit product cost

Product S Product W
Machine setup 157300 323700
Purchase order 42354 10686
Machine hour 116700 169650
Maintenance request 93408 125100
Total overhead 409762 629136
Unit 70500 14100
Overhead cost per unit 5.81 44.62
Direct material 14 29
Direct labor 15 14
Unit product cost 34.81 87.62